# Google Research

### v1

@@ Appendix C Summary Comparison Examples
Appendix D Additional Related Work
@@
Our work builds on real-world conversations with LLM chatbots from recent open-source datasets (Zheng et al., 2023; Zhao et al., 2024). Prior work has analyzed such conversations to extract structure and insight. For instance, Zhao et al. (2024) define a set of common conversation categories and prompt LLMs to classify conversations accordingly. In contrast, Zheng et al. (2023) and Tamkin et al. (2024) adopt more flexible approaches that do not rely on predefined categories. Zheng et al. (2023) cluster conversations in the embedding space and then use LLMs to summarize each cluster’s main theme. Tamkin et al. (2024) build on this by first summarizing individual conversations into concise summaries and then applying hierarchical clustering. While these works provide valuable frameworks for understanding LLM behavior, none offer formal privacy guarantees. Our work addresses this gap by introducing the first differentially private conversation summarization pipeline.
@@
Both Tamkin et al. (2024) and our work employ a clustering step to group conversations with similar embeddings together. We use $k$-means, which is one of the most popular formulations of clustering; see e.g. Ikotun et al. (2023) for a comprehensive survey. For non-DP $k$-means, we use the classic algorithm of Lloyd (1982). For DP $k$-means, several algorithms have been proposed, starting with Blum et al. (2005) who devised a DP version of Lloyd’s algorithm. This was followed up by a series of works improving different algorithmic guarantees (e.g. Nissim et al. (2007); Gupta et al. (2010); Balcan et al. (2017); Ghazi et al. (2020)), culminating in the algorithm of Chang et al. (2021) which not only gives a nearly optimal (theoretical) approximation ratio but is also practical. We ended up using the open-source library Chang & Kamath (2021) based on this paper. We remark that there are also more recent works which provide practical DP clustering algorithms, e.g.  Tsfadia et al. (2022); nevertheless, we are unaware of any open-source library based on these papers.
@@
Our empirical privacy evaluation is in spirit similar to the MIA. MIA aims to infer whether specific data points were included in the training data of a given model or algorithm, based on its outputs. While various MIA techniques have been proposed, they primarily target machine learning models (Shokri et al., 2017; Yeom et al., 2018; Sablayrolles et al., 2019; Carlini et al., 2022). In this work, we design an empirical privacy evaluation method tailored to attack the LLM conversation summarization pipeline.
@@
Our keyword set generation method (section 4.1) is related to recent work on extracting keywords from text corpora using LLMs (Grootendorst, 2020; Maragheh et al., 2023; Wang et al., 2024). These approaches typically fine-tune or prompt an LLM to extract keywords. However, such pipelines are not applicable in a DP setting, as documents may contain unique keywords that could leak sensitive information. In this work, we propose keyword generation methods that only output keywords shared across multiple conversations, allowing us to provide formal DP guarantees.
Appendix E Empirical Privacy Leakage Evaluation
@@ Appendix F LLM Prompts
F.1 Extracting Embeddings from Records
Below, we provide the prompt used in the $\mathsf{ExtractEmbeddings}$ method in Algorithm 4 and Algorithm 1. The input record $x$ is substituted in place of $\langle x\rangle$ in the prompt. Note that we apply the embedding function specifically to the summary portion of the LLM response obtained using this prompt, rather than to the entire response.
$\mathsf{ExtractEmbeddings}(x)$adjustwidth} -4em-4em
@@ F.1 Extracting Embeddings from Records
F.2 LLM-Summarize in Algorithm 4
We provide the prompt used for LLM-Summarize method in Algorithm 4. In both the Clio and our proposed pipeline implementations, we occasionally select contrastive summaries from the nearest points located outside the designated cluster. For the sake of algorithmic clarity, we omitted this nuanced selection strategy from Algorithm 4 to maintain a simplified representation of the core approach.
LLM-Summarize {adjustwidth} -4em-4em
@@ F.4 LLM-Summarize in Algorithm 1
F.5 Prompt for KwSet-LLM
@@
KwSet-LLM Generation Before Refinement {adjustwidth} -4em-4em
@@
As a KwSet of smaller size leads to smaller utility loss due to privacy, we asked LLM to refine the KwSet set.
Refine KwSet {adjustwidth} -4em-4em
@@
F.6 Prompt for Evaluation
@@
Prompt for Binary Preference {adjustwidth} -4em-4em
@@
Appendix G Hierarchy and Visualization
G.1 Hierarchical Organization
@@
    Low-level Summary Generation: We first generate approximately 4,000 low-level summaries using our differentially private pipeline.
@@
This hierarchical organization allows users to navigate from broad topics to specific conversation summaries, making the system more useful for exploring large conversation datasets while maintaining differential privacy guarantees. Some examples of the summary facet, base, and top cluster are demonstrated in Table 4.
G.2 Visualization Interface
To improve the utility and exploration capabilities of our hierarchical organization, we developed interactive visualization interfaces that allow users to navigate through both high-level and low-level clusters. Our visualization system offers two complementary views:
Refer to caption Figure 5: * (a) Tree View Refer to caption Figure 6: * (b) UMAP-based Map View
Refer to caption Figure 7: * (c) Detailed view of a selected cluster
Figure 8: Visualization interfaces for exploring hierarchical clusters: (a) Tree View displaying the hierarchical organization with nested clusters and their sizes; (b) Map View showing semantic relationships between topics through UMAP dimensionality reduction, with similar topics positioned closer together; (c) Detailed view of a selected cluster showing its summary and constituent sub-clusters with their specific focus areas.
Tree View.
The tree visualization (Figure 8a) presents clusters in a hierarchical format, with high-level clusters containing their associated low-level clusters. One can click the button for each high-level cluster to show and hide the associated low-level clusters. This view enables users to navigate from broad topics (e.g., ”AI Ethics, Harmful Content, and Societal Impact”) to more specific subtopics (e.g., ”AI Ethics, Society, and Cultural Impact”). The tree structure includes the number of contained conversations for each cluster, providing a sense of topic prevalence.
Map View.
The map visualization (Figure 8b) projects cluster embeddings into a two-dimensional space using UMAP McInnes et al. (2018). This representation preserves semantic relationships between clusters, with similar topics appearing closer together. The map view offers an intuitive understanding of the thematic landscape and reveals natural groupings of related topics. Users can toggle between high-level and low-level views.
When a user selects a specific cluster (Figure 8c), the interface displays detailed information about that cluster, including its summary and a list of its sub-clusters. For high-level clusters, users can explore the constituent low-level clusters, each with its own descriptive title. This allows users to quickly understand the scope and content of each cluster without needing to examine individual conversations.
This visualization system enhances the utility of our differentially private summarization framework by making the results more navigable and interpretable, even when dealing with thousands of conversation summaries organized into a complex hierarchy.

### v2

@@ Appendix D Summary Comparison Examples
Appendix E Additional Related Work
@@
Our work builds on real-world conversations with LLM chatbots from recent open-source datasets (Zheng et al., 2023; Zhao et al., 2024). Prior work has analyzed such conversations to extract structure and insight. For instance, Zhao et al. (2024) define a set of common conversation categories and prompt LLMs to classify conversations accordingly. In contrast, Zheng et al. (2023) and Tamkin et al. (2024) adopt more flexible approaches that do not rely on predefined categories. Zheng et al. (2023) cluster conversations in the embedding space and then use LLMs to summarize each cluster’s main theme. Tamkin et al. (2024) build on this by first summarizing individual conversations into concise summaries and then applying hierarchical clustering. While these works provide valuable frameworks for understanding LLM behavior, none offer formal privacy guarantees. Our work addresses this gap by introducing the first DP conversation summarization pipeline.
@@
Both Tamkin et al. (2024) and our work employ a clustering step to group conversations with similar embeddings together. We use $k$-means, which is one of the most popular formulations of clustering; see e.g. Ikotun et al. (2023) for a comprehensive survey. For non-DP $k$-means, we use the classic algorithm of Lloyd (1982). For DP $k$-means, several algorithms have been proposed, starting with Blum et al. (2005) who devised a DP version of Lloyd’s algorithm. This was followed up by a series of works improving different algorithmic guarantees (e.g. Nissim et al. (2007); Gupta et al. (2010); Balcan et al. (2017); Ghazi et al. (2020)), culminating in the algorithm of Chang et al. (2021) which not only gives a nearly optimal (theoretical) approximation ratio but is also practical. We ended up using the open-source library Chang & Kamath (2021) based on this paper. We remark that there are also more recent works which provide practical DP clustering algorithms, e.g.,  Tsfadia et al. (2022); nevertheless, we are unaware of any open-source library based on these papers.
@@
Our empirical privacy evaluation (§F) is in spirit similar to MIA, which aims to infer whether specific data points were included in the training data of a given model or algorithm, based on its outputs. While various MIA techniques have been proposed, they primarily target machine learning models (Shokri et al., 2017; Yeom et al., 2018; Sablayrolles et al., 2019; Carlini et al., 2022). In this work, we design an empirical privacy evaluation method tailored to attack the LLM conversation summarization pipeline.
@@
Our keyword set generation method (§4.1) is related to recent work on extracting keywords from text corpora using LLMs (Grootendorst, 2020; Maragheh et al., 2023; Wang et al., 2024). These approaches typically fine-tune or prompt an LLM to extract keywords. However, such pipelines are not applicable in a DP setting, as documents may contain unique keywords that could leak sensitive information. In this work, we propose keyword generation methods that only output keywords shared across multiple conversations, allowing us to provide formal DP guarantees.
Appendix F Empirical Privacy Leakage Evaluation
@@ Appendix G LLM Prompts
G.1 Extracting Embeddings from Records
Below, we provide the prompt used in the $\mathsf{ExtractEmbeddings}$ method in Algorithm 4 and Algorithm 1. The input record $x$ is substituted in place of $\langle x\rangle$ in the prompt. Note that we apply the embedding function specifically to the ‘Summary’ portion of the LLM response obtained using this prompt, rather than to the entire response.
Prompt for $\mathsf{ExtractEmbeddings}(x)$
@@ G.1 Extracting Embeddings from Records
G.2 LLM-Summarize in Simple-Clio (Algorithm 4)
We provide the prompt used for LLM-Summarize method in Simple-Clio. In both Clio and our proposed pipeline implementations, we select contrastive summaries from the nearest points located outside the designated cluster. For the sake of algorithmic clarity, we omitted this nuanced selection strategy from Algorithm 4 to maintain a simplified representation of the core approach.
LLM-Summarize
@@ G.4 LLM-Summarize in Algorithm 1
G.5 Prompts for KwSet-LLM and KwSet-TFIDF
@@
KwSet-LLM Generation (Before Refinement)
@@
As a keyword set of smaller size leads to smaller utility loss due to privacy, we use an LLM to refine the keyword set (for both KwSet-TFIDF and KwSet-LLM) using the following prompt template.
Refine KwSet
@@
G.6 Prompt for KwSet-Public
Refine KwSet-Public
You are refining a set of extracted key phrases for topic summarization.
Below is a list of candidate keywords extracted from conversations:
[‘, ’.join(all_candidate_keywords)]
Your task:
- **Remove meaningless words** such as random numbers, dates, single letters, generic words (e.g., "thing", "something", "stuff").
- **Delete overly specific entity names** (e.g., "John Doe", "xyz123", "April 5, 2023", "ID number").
- **Remove redundant words** that mean the same thing.
- **Merge similar concepts** (e.g., "AI fairness" and "bias in AI" should be unified).
- **Keep only the most informative, meaningful terms** that are useful for summarizing topics.
- **Avoid vague or overly broad words** like "technology", "question", "people".
- **Ensure the final keywords are suitable for clustering and summarization**.
Return only the **refined keyword set** as a comma-separated list.
Update KwSet-Public by adding and removing keywords
You are an expert in topic summarization and keyword analysis. The current public keyword set contains {len(existing_keywords)} words:
[{‘, ’.join(existing_keywords))}]
Here is a batch of private conversations:
‘{batch_text}’
Your task:
- Compare the batch of private conversations with the current keyword set.
- Identify keywords in the current set that are redundant or irrelevant.
- Identify new, meaningful keywords present in the private conversations that are not in the current set.
- Output ONLY a complete, valid JSON object (with no extra text) with the following structure:
   ‘words_to_remove’: [list of words to remove],
   ‘words_to_append’: [list of words to add]
Ensure your output starts with ‘{’ and ends with ‘}’ (i.e., the JSON object must be complete, including the closing bracket). Keep your answer as concise as possible.
G.7 Prompt for KwSet-Hybrid
Refine KwSet-Hybrid
You are an expert in topic analysis and keyword extraction.
The current public keyword set contains {len(public_keywords)} words:
[{‘, ’.join(public_keywords)}]
Here is a batch of private conversations:
‘{batch_text}’
Your task:
- Analyze the above conversations and identify any new, meaningful keywords that are not already present in the public set.
- Omit any non-English conversations or keywords.
- Return ONLY a concise, comma-separated list of candidate keywords with no additional explanation of at most 5 words.
G.8 Prompt for Evaluation
@@
Prompt for Binary Preference
@@
Appendix H Hierarchy and Visualization
H.1 Hierarchical Organization
@@
    Low-level Summary Generation: We first generate approximately 4,000 low-level summaries using our DP pipeline.
@@
This hierarchical organization allows users to navigate from broad topics to specific conversation summaries, making the system more useful for exploring large conversation datasets while maintaining DP guarantees. Some examples of the summary facet, base, and top cluster are demonstrated in Table 4.